# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v2

@@ 4.3.1 Effect of CoT Prompting
We examine how CoT prompting improves LLM robustness through explicit step-by-step execution. From Table 3, we observe that CoT effectively enhances model robustness, with relative improvements of $+10.01$pp for ALL, $+9.23$pp for MDC, $+4.04$pp for MPS, and $+13.57$pp for MHC. The results indicate that step-by-step reasoning guides LLMs to prioritize executable semantics over irrelevant code segments and textual cues. Notably, the substantial improvement in MHC suggests CoT effectively mitigates the over-reliance on reasoning-level hints by re-centering reasoning around executable semantics. Such benefits are particularly evident in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), reflecting enhanced reasoning capabilities and greater noise resilience.
However, CoT prompting cannot entirely eliminate the model’s sensitivity to misleading NL cues and structural distractions. Substantial performance degradation remains under ALL, MDC, and MPS perturbations, particularly for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1). These results underscore an intrinsic limitation of current models in distinguishing non-functional code and misleading information from essential logic. Interestingly, models usually incorporate print statements directly into their reasoning paths, leading to greater degradation under MPS compared to MDC due to self-induced misinterpretation. Since CoT prompting exposes intermediate reasoning processes, we provide detailed case studies for each perturbation in Appendix E (Case 1 to 4) to investigate model inference and illustrate their failure reasons.
Key Finding: CoT prompting significantly improves LLM robustness by enforcing explicit step-by-step reasoning, particularly benefiting larger models. Nevertheless, it cannot fully eliminate models’ reliance on misleading cues, indicating fundamental limitations in reasoning capabilities.
4.3.2 Powerful Reasoning and Potential Vulnerability of LRMs
We then evaluate the reliability of three LRMs on output prediction tasks, with results summarized in Table 4. Compared to LLMs, LRMs show outstanding performance and robustness, surpassing $90$ Pass@1 in vanilla settings and maintaining minimal performance drop under ALL and MPS perturbations. These results indicate that internal reasoning outperforms CoT prompting in accurately distinguishing essential executable code segments from irrelevant code and misleading distractions.
Notably, LRMs consistently exhibit stronger resistance to MPS compared to MDC, suggesting an inherent cognitive bias toward treating comments as authoritative execution guidance while appropriately disregarding print content. However, this residual cognitive bias exposes LRMs to critical vulnerabilities under reasoning-level perturbation that plausible but incorrect hints can trigger severe reasoning failures. Specifically, we observe the following outcomes under MHC: (i) substantial increases in reasoning token usage across all LRMs, often doubling when compared to VAN; (ii) extreme cases in QwQ-32B—occurring in $4/479$ tasks; see Appendix E.6—where it generates over 32,000 tokens due to uncontrolled recursive self-verification and reasoning collapse; (iii) increased reasoning effort and token usage instead exacerbate performance drops in o3-Mini, suggesting that internal reasoning can amplify confusion when led by contradiction. These findings reveal that reasoning-level contradictories can evoke severe internal reasoning instability, resulting in significant computational overhead, excessive self-reflection, and even severe performance drops.
@@ 4.3.2 Powerful Reasoning and Potential Vulnerability of LRMs
Table 4: Absolute Pass@1 degradation and reasoning token usage (Avg. and Max.) of LRMs on output prediction across Crux and LCB ($N=1$).
Model Dataset VAN ALL MDC MPS MHC
@@
o3-Mini-Low Crux 97.62 $213$ 2560 $-1.75$ $379$ 2240 $-4.25$ $244$ 5184 $-1.00$ $272$ 3584 $-10.88$ $366$ 4864
LCB 98.96 $240$ 1024 $-0.63$ $543$ 1856 $-12.32$ $265$ 1088 $-3.55$ $280$ 1088 $-19.62$ $330$ 2752
o3-Mini-High Crux 98.12 $1311$ 20000 $+0.12$ $2223$ 20000 $-3.50$ $2182$ 20000 $+0.88$ $2197$ 20000 $-13.12$ $3108$ 20000
LCB 100.00 $1084$ 8576 $-0.21$ $2632$ 8960 $-5.64$ $2136$ 8448 $-0.63$ $1972$ 7744 $-28.39$ $3767$ 20000
DeepSeek-R1 Crux 95.38 $922$ 10471 $-1.25$ $1469$ 11191 $-3.38$ $1424$ 10807 $-0.38$ $1069$ 10037 $-2.25$ $2216$ 16021
LCB 99.79 $906$ 7322 $-1.25$ $1618$ 7749 $-2.71$ $1512$ 6157 $-0.42$ $1094$ 9378 $-0.63$ $2597$ 14870
QwQ-32B Crux 93.25 $1391$ 13718 $-0.88$ $2088$ 12180 $-3.12$ $1814$ 10906 $-1.12$ $1875$ 11700 $-0.75$ $2657$ 18759
LCB 98.96 $1518$ 9206 $-0.21$ $2488$ 11232 $-4.59$ $1669$ 8991 $-1.88$ $1751$ 8816 $-1.04$ $3713$ 32769
Key Finding: LRMs leverage strong internal reasoning capabilities to achieve remarkable performance and robustness. However, their cognitive tendency toward comments introduces severe vulnerabilities, resulting in doubled token usage and reasoning collapse under MHC perturbation.
4.4 Impact on Input Prediction
Table 5: Comparison of relative Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
Model Series Model Name VAN ALL MDC MPS MHC
@@ 4.4 Impact on Input Prediction
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-29.84\%$ $-19.63\%$ $+1.30\%$ $+1.05\%$ $+1.73\%$ $+1.05\%$ $-30.49\%$ $-8.66\%$
GPT-4o $68.04$ $79.12$ $-13.29\%$ $-7.42\%$ $+5.63\%$ $+1.26\%$ $+1.41\%$ $+2.69\%$ $-8.33\%$ $-2.84\%$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-17.00\%$ $-16.70\%$ $+7.01\%$ $+4.13\%$ $+7.77\%$ $+2.36\%$ $-45.35\%$ $-29.86\%$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-15.91\%$ $-9.67\%$ $+0.11\%$ $-3.43\%$ $-1.70\%$ $-2.65\%$ $-24.83\%$ $-17.00\%$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-30.15\%$ $-23.77\%$ $+10.20\%$ $-5.43\%$ $+6.00\%$ $-4.24\%$ $-30.95\%$ $-11.04\%$
Gemini-2.0-Flash $70.50$ $84.88$ $-41.31\%$ $-13.40\%$ $-0.77\%$ $-2.80\%$ $-11.47\%$ $-12.08\%$ $-21.51\%$ $-4.27\%$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-20.87\%$ $-14.96\%$ $-2.47\%$ $-7.02\%$ $-3.76\%$ $-9.62\%$ $-18.69\%$ $-7.63\%$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-16.59\%$ $-7.99\%$ $+1.12\%$ $-0.30\%$ $+1.57\%$ $+0.90\%$ $-5.61\%$ $+3.17\%$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-30.53\%$ $-61.86\%$ $+3.87\%$ $+1.50\%$ $+3.03\%$ $-3.60\%$ $-7.18\%$ $-24.92\%$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-28.29\%$ $-17.89\%$ $+4.91\%$ $+4.52\%$ $+1.61\%$ $+1.69\%$ $-17.41\%$ $-10.17\%$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-15.28\%$ $-9.00\%$ $+5.86\%$ $+3.67\%$ $+2.99\%$ $+2.33\%$ $-7.80\%$ $-5.67\%$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-29.39\%$ $-15.69\%$ $+1.25\%$ $-0.33\%$ $-4.08\%$ $+0.65\%$ $-21.83\%$ $-11.11\%$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-44.42\%$ $-50.36\%$ $+14.70\%$ $+3.16\%$ $+12.23\%$ $+7.54\%$ $-9.33\%$ $-14.36\%$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-38.42\%$ $-15.02\%$ $+8.66\%$ $+10.91\%$ $+4.86\%$ $+5.14\%$ $-8.74\%$ $-1.23\%$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-26.05\%$ $-20.40\%$ $+7.74\%$ $+0.51\%$ $+2.23\%$ $-4.05\%$ $-20.60\%$ $-10.96\%$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-26.86\%$ $-9.88\%$ $+2.65\%$ $+1.61\%$ $-3.68\%$ $+2.46\%$ $-25.53\%$ $-6.86\%$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-28.56\%$ $-20.61\%$ $-1.01\%$ 0.00 $-2.41\%$ $+0.32\%$ $-36.48\%$ $-33.39\%$
Average $\mathbf{-26.63}\%$ $\mathbf{-19.66}\%$ $\mathbf{+4.16}\%$ $\mathbf{+0.77}\%$ $\mathbf{+1.08}\%$ $\mathbf{-0.54}\%$ $\mathbf{-20.04}\%$ $\mathbf{-11.58}\%$
@@
Finally, we analyze how perturbations affect LLMs in input prediction; the results are summarized in Table 5. Consistent with our output prediction findings in §4.2, ALL and MHC perturbations substantially degrade model performance, causing relative average drops of $26.63\%$ and $20.04\%$, respectively. These results support Gu et al. (2024)’s hypothesis that robustness in related coding tasks is interrelated. CoT prompting also improves model performance by $+6.97$pp under ALL and $+8.64$pp under MHC. However, improvements in input prediction are less effective than in output prediction due to the unique challenges of the inverse nature of the reasoning process.
@@
Surprisingly, in contrast to output prediction, contextual-level textual perturbations (MDC and MPS) do not negatively impact LLMs in input prediction tasks. Since input prediction tasks accept multiple valid solutions, models can bypass reasoning about certain code branches for a correct answer. Our misleading NL messages, initially designed to confuse and distract models, inadvertently reinforce such shortcuts, simplifying models’ reasoning processes. Consequently, these perturbations unintentionally provide shallow reasoning strategies, so those improvements do not reflect the genuine robustness of LLMs in code comprehension. We provide a detailed case study in Appendix E.5, illustrating how models adopt misleading messages to shortcut their reasoning.
Key Finding: ALL and MHC significantly degrade performance on input prediction, highlighting fundamental reasoning limitations. Conversely, misleading messages accidentally simplify reasoning, prompting models to produce superficial but valid answers without true understanding.
@@ 5 Discussion
5.1 Comment Density and Robustness Degradation
Figure 1: Relative Pass@1 degradation under varying MDC injection probabilities $p$.
@@ 5.1 Comment Density and Robustness Degradation
Table 6: Reasoning token usage under varying MDC injection probabilities $p$.
We design a textual density stress test to confirm the validity of MDC perturbations and evaluate model robustness under varying distraction densities on Crux dataset. Specifically, we examine how relative Pass@1 drop across different Qwen models as we gradually increase the comment injection probability $p\in\{0.0,0.1,\dots,1.0\}$, illustrated in Figure 1. We further analyze the role of reasoning by evaluating CoT inference using Qwen2.5-72B and internal reasoning using QwQ-32B, with detailed results on token usage summarized in Table 6.
All models exhibit a clear performance degradation trend as injection probability $p$ increases, despite minor non-linearities due to randomness in insertion positions and comment content. Notably, smaller models degrade more rapidly, while larger models demonstrate greater resilience at high densities. This trend persists in reasoning-enhanced models (CoT and LRMs), showing superior robustness against dense textual distractions. These results confirm the validity of injecting misleading messages (MDC and MPS) as a scalable stress-testing technique and an effective strategy to evaluate LLM robustness under contextual-level perturbations.
While Qwen2.5-72B with CoT prompting and QwQ-32B have similar trends in the stress test, they respond and reason differently under dense comment noise. CoT prompting improves model robustness by guiding models to selectively focus on the essential code logic and ignore distracting messages. However, it can lead to premature conclusions without sufficiently analyzing the code, so token usage is slightly reduced as models do not analyze the comments and reduce the inference effort. In contrast, QwQ-32B processes all available information and attempts an exhaustive analysis, actively resolving contradictions before making a final decision. While this comprehensive analysis achieves higher vanilla performance and overall robustness, it risks severe reasoning collapse under heavily contradictory conditions, as discussed in §4.3.2.
5.2 Prompting to Ignore Misleading Comments

### v3

@@ 3.3 Limits of CoT in Resolving NL-Embedded Contradictions
CoT encourages pre-decision deliberation, which enhances resilience to irrelevant segments and embedded cues, reducing relative performance degradation from $23.2\%$ to $13.8\%$ on average. Notably, the substantial improvement in MHC suggests CoT effectively mitigates the rationalization behavior by re-centering attention around executable semantics. Such benefits are particularly evident in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), reflecting enhanced code-tracing ability, stronger critical reasoning capabilities, and greater noise resilience. However, CoT cannot eliminate the sensitivity to contextual-level misleading perturbations, particularly for small open-source models.
Failure Case Analysis: Inconsistency and Rationalization in CoT.
From Table 3, we observe clear failure modes showing how Qwen2.5-72B is misled. Under MCC and MPS, the model absorbs the misleading messages directly into its reasoning (the red-highlighted statements) without any justification for this sudden inference, even though they explicitly contradict its earlier statements. Under MHC, the model initially conducts correct reasoning for the fourth character ‘w’, but in the very next step, ‘a’, it produces an incorrect conclusion to force alignment with the misleading hint, contradicting its previous inference. These behaviors align with the distractibility Shi et al. (2023) and post-hoc rationalization Barez et al. (2025), where the model overrides its correct reasoning to conform to injected context. These issues are particularly severe in programming because comments and prints are semantically irrelevant to execution, but models possibly treat them as higher-priority evidence than program logic, which underscores an intrinsic limitation of LLMs in distinguishing non-functional code and misleading information from essential logic, highlighting insufficient critical reasoning capability. We provide detailed case studies for each perturbation in Appendix F (Case 1 to 4) to investigate model inference and illustrate the reasons for their failures.
3.4 Powerful Reasoning and Instability of LRMs
@@
Table 4: Relative Pass@1 degradation and reasoning token usage (Avg. and Max.) of LRMs for output prediction across Crux and LCB ($N=1$).
Model Dataset VAN PSC-ALL MCC MPS MHC
@@
o3-Mini-Low Crux $97.6$ $213$ $2560$ $1.8\%$ $379$ $2240$ $4.4\%$ $244$ $5184$ $1.0\%$ $272$ $3584$ $11.1\%$ $366$ $4864$
LCB $99.0$ $240$ $1024$ $0.6\%$ $543$ $1856$ $12.4\%$ $265$ $1088$ $3.6\%$ $280$ $1088$ $19.8\%$ $330$ $2752$
o3-Mini-High Crux $98.1$ $1311$ $20000$ $0.1\%$ $2223$ $20000$ $3.6\%$ $2182$ $20000$ $0.9\%$ $2197$ $20000$ $13.4\%$ $3108$ $20000$
LCB $100$ $1084$ $8576$ $0.2\%$ $2632$ $8960$ $5.6\%$ $2136$ $8448$ $0.6\%$ $1972$ $7744$ $28.4\%$ $3767$ $20000$
DeepSeek-R1 Crux $95.4$ $929$ $10542$ $1.3\%$ $1477$ $11101$ $3.5\%$ $1436$ $10927$ $0.4\%$ $1078$ $10150$ $4\% 22$332233 $16079$
LCB $99.8$ $909$ $7347$ $1.3\%$ $1621$ $7759$ $2.7\%$ $1519$ $6187$ $0.4\%$ $1099$ $9398$ $0.6\%$ $2605$ $14889$
QwQ-32B Crux $93.2$ $1409$ $14263$ $0.9\%$ $2110$ $12499$ $3.4\%$ $1834$ $10959$ $1.2\%$ $1895$ $11935$ $0.8\%$ $2694$ $19491$
LCB $99.0$ $1530$ $9230$ $0.2\%$ $2517$ $11232$ $4.6\%$ $1681$ $8993$ $1.9\%$ $1763$ $8818$ $1.1\%$ $3740$ $32764$
Figure 2: Quadratic increase ($R^{2}=0.9991$) in confusion tokens (e.g., “hmm”, “wait”, “perhaps”) during QwQ-32B’s reasoning process.
Figure 3: Breakdown of QwQ-32B’s reasoning trajectory under MHC perturbation on LCB sample 259.
Building on the findings in §3.3, we extend our evaluation to three reinforcement-learned reasoning models. As shown in Table 4, LRMs demonstrate outstanding performance and robustness under perturbations, with negligible degradation under PSC-ALL, outperforming CoT-prompted LLMs in code tracing, logic understanding, and segment distinguishing. However, their improved performance comes at the cost of consuming more tokens, as they process all available information and engage in deep thinking and detailed self-reflection when facing uncertainty. Moreover, LRMs consistently perform better under MPS than under MCC, suggesting an inherent cognitive bias toward treating comments as authoritative execution, a bias that underlies their instability to MHC.
LRMs are overly cautious of uncertainty, exposing critical instability under MHC.
Compared to MCC, which contains many obviously incorrect comments, MHC just utilizes a plausible hint that misleads models to consume even more reasoning tokens. Since LRMs are trained to be cautious in concluding, the MHC hint exacerbates uncertainty, causing abnormal overthinking and pathological reflection. As shown in Table 4, this manifests in three notable outcomes: (i) reasoning token usage increases by $2-3\times$ compared to the vanilla setting, (ii) in extreme cases, such as QwQ-32B, it generates over 32k tokens due to uncontrolled recursive self-verification, and (iii) for o3-Mini, increased reasoning effort further exacerbates performance degradation, indicating that internal reasoning can amplify rather than resolve uncertainty. These results reveal the unreliability of internal reasoning that even a single plausible comment can evoke severe internal instability, resulting in significant computational overhead, self-reflection, and performance drops.
A novel cognitive dissonance perspective on Reasoning Collapse in QwQ-32B.
For QwQ-32B, we observed four “collapse” events (i.e., generating massive replicated terms), all from MHC on LCB, suggesting that this is not a coincidental phenomenon. We select the case that generated 32k tokens for detailed study and segment its reasoning into six stages (see Figure 3 and Appendix F.6). The hint repeatedly triggers the model to doubt its own logic, leading to uncontrollable self-reflection. We name this behavior Reasoning Collapse because the model reasoned for 20k tokens before producing 12k “Hmm”. Figure 3 shows a quadratic growth of confusing tokens before the collapse, indicating that the failure is attributable to the accumulated conflict and growing uncertainty rather than a low-level glitch. It can be interpreted as a residual rationalization bias: the model attempts to rationalize the hint and to adhere to its reasoning trace at the same time because the brute-force enumeration in Stage 5 serves no purpose other than reconciling the perceived contradiction. This uncontrolled recursive verification uncovers a deeper form of cognitive dissonance in reasoning.
3.5 Potential Limitation in Input Prediction
Table 5: Relative Pass@1 degradation for input prediction tasks across different reasoning modes on Crux. See [Absolute Version] for the corresponding absolute degradation.
Model Series Model Name VAN PSC-ALL MCC MPS MHC
@@ 3.5 Potential Limitation in Input Prediction
GPT-4 Omni GPT-4o-Mini $57.8$ $69.5$ $29.9\%$ $19.6\%$ $1.2\% +$$1.0\% +$$1.4\% +$$1.0\%$ $30.6\%$ $8.7\%$
GPT-4o $68.0$ $79.1$ $13.3\%$ $7.4\%$ $5.6\% +$$1.3\% +$$1.4\% +$$2.7\%$ $8.3\%$ $2.8\%$
Claude Series Claude-3.5-Haiku-20241022 $54.7$ $63.6$ $17.1\%$ $16.7\%$ $6.8\% +$$4.1\% +$$7.8\% +$$2.4\%$ $45.4\%$ $29.9\%$
Claude-3.5-Sonnet-20241022 $73.3$ $80.1$ $15.9\%$ $9.7\%$ $0.1\%$ $3.4\%$ $1.8\%$ $2.7\%$ $24.8\%$ $17.0\%$
Gemini Series Gemini-1.5-Flash $57.6$ $73.6$ $29.9\%$ $23.8\%$ $10.2\%$ $5.4\%$ $6.0\%$ $4.2\%$ $30.9\%$ $11.0\%$
Gemini-2.0-Flash $70.5$ $84.9$ $42.1\%$ $13.4\%$ $0.8\%$ $2.8\%$ $11.5\%$ $12.1\%$ $21.5\%$ $4.3\%$
Gemini-1.5-Pro-002 $71.0$ $81.9$ $21.0\%$ $14.8\%$ $2.5\%$ $7.0\%$ $3.8\%$ $9.6\%$ $18.8\%$ $7.6\%$
DeepSeek DeepSeek-V3 $69.1$ $82.9$ $16.6\%$ $8.0\%$ $1.1\%$ $0.3\%$ $1.7\% +$$0.9\%$ $5.9\%$ $3.2\%$
LLaMA Series LLaMA-3.1-8B-Instruct $37.1$ $41.6$ $30.8\%$ $61.9\%$ $3.6\% +$$1.5\% +$$3.0\%$ $3.6\%$ $7.2\%$ $24.9\%$
LLaMA-3.1-70B-Instruct $62.1$ $66.4$ $28.4\%$ $17.7\%$ $4.7\% +$$4.5\% +$$1.6\% +$$1.9\%$ $17.4\%$ $10.0\%$
LLaMA-3.1-405B-Instruct $66.8$ $75.0$ $15.2\%$ $9.0\%$ $5.9\% +$$3.7\% +$$2.9\% +$$2.5\%$ $7.9\%$ $5.7\%$
LLaMA-3.3-70B-Instruct $63.3$ $76.5$ $29.1\%$ $15.5\%$ $1.2\%$ $0.3\%$ $4.0\%$ $0.7\%$ $21.8\%$ $11.1\%$
Qwen Series Qwen2.5-7B-Instruct $38.8$ $51.4$ $44.4\%$ $50.4\%$ $14.7\% +$$3.4\% +$$12.2\% +$$7.5\%$ $9.4\%$ $14.4\%$
Qwen2.5-14B-Instruct $50.4$ $60.8$ $38.4\%$ $15.0\%$ $8.4\% +$$10.9\% +$$\% +5.3+5.3\%$ $8.6\%$ $1.2\%$
Qwen2.5-32B-Instruct $63.5$ $74.1$ $26.0\%$ $20.4\%$ $7.4\% +$$0.5\% +$$2.2\%$ $4.0\%$ $20.6\%$ $11.0\%$
Qwen2.5-72B-Instruct $64.6$ $74.1$ $27.0\%$ $10.8\%$ $2.6\% +$$1.2\%$ $3.7\%$ $2.0\%$ $25.5\%$ $7.3\%$
Qwen2.5-Coder-32B-Instruct $74.2$ $78.4$ $28.6\%$ $20.6\%$ $1.0\%$ $0.2\%$ $2.4\%$ $0.2\%$ $6.5\% 33$$3.5\%$
Average $\mathbf{-26.7\%}$ $\mathbf{-19.6\%}$ $\mathbf{+4.1\%}$ $\mathbf{+0.8\%}$ $\mathbf{+1.1\%}$ $\mathbf{-0.5\%}$ $\mathbf{-20.1\%}$ $\mathbf{-11.5\%}$
@@
We repeat the experiments in input prediction tasks on the Crux dataset and summarize the results in Table 5. The results under PSC-ALL and MHC align with our findings in §3.2 and §3.3, further supporting Gu et al. (2024)’s hypothesis that robustness in related coding tasks is interrelated. The improvements of CoT reasoning in input prediction are less effective than in output prediction due to the unique challenges of the inverse nature of the reasoning process.
Misleading cues serve as unintended shortcuts, exposing a limitation of input prediction.
@@
Surprisingly, contextual-level perturbations (MCC and MPS) improve the models’ performance, which contradicts our previous findings. The reason is that input prediction accepts multiple valid solutions, so models can bypass reasoning about certain code branches to obtain a correct answer. Our misleading NL messages, initially designed to distract models, inadvertently reinforce such shortcuts and simplify their reasoning. Consequently, these perturbations unintentionally provide shallow reasoning strategies, so those improvements do not reflect the genuine robustness of LLMs in code comprehension. We provide a detailed case study in Appendix F.5, illustrating how models adopt misleading messages to shortcut their reasoning.
@@ 4 Discussion
4.1 Impact of Comment Density on Robustness and Reasoning Behaviors
Figure 4: Relative Pass@1 degradation under the varying MCC injection probabilities pp.
Table 6: Reasoning token usage under the varying MCC injection probabilities pp.
@@ 4.1 Impact of Comment Density on Robustness and Reasoning Behaviors
To confirm the validity and realism of our misleading perturbations, we design a density sweep experiment to examine how the relative Pass@1 score drops across different models when we gradually increase the injection probability $p\in\{0.0,0.1,\dots,1.0\}$, as illustrated in Figure 4.
Systematic degradation under increasing density.
Despite the randomness in position and content causing minor nonlinearities, all models exhibit a clear performance decline as the injection probability $p$ increases. Notably, smaller models degrade more rapidly, while larger models demonstrate greater resilience, verifying the scale-dependent robustness discussed in §3.2. These results confirm that contextual-level misleading perturbations are a scalable stress-testing method. Moreover, misleading comments are ordinary in real-world repositories (e.g., CPython GitHub issue  [^3] See discussion in https://github.com/python/cpython/issues/136764.), and our results verify that even a single misleading comment can impact model understanding.
LLMs and LRMs have similar trends but different reasoning behaviors.
Qwen2.5-72B (CoT) and QwQ-32B have similar trends in performance degradation, but they have opposite trends in token usage, as shown in Table 6. While CoT improves robustness by enabling step-by-step execution, LLMs solely report the expected behavior and variable changes (as shown in Table 3), causing them to reach premature conclusions without sufficiently analyzing the code. Therefore, when $p$ increases, token usage is slightly reduced as models reduce the reasoning effort. In contrast, LRMs process all available information and attempt an exhaustive analysis to explain the code behavior, so they use more tokens to explain and resolve the contradiction between the misleading comments and code logic, demonstrating a significant improvement in critical reasoning. This critical and comprehensive analysis achieves higher performance and robustness, but it risks severe reasoning collapse if it cannot resolve the contradiction as discussed in §3.4.
4.2 Cross-Language Robustness Analysis
Table 7: Performance drop across Python (PY) and JavaScript (JS) under different perturbations.
Model Inference VAN PSC-ALL MCC MPS MHC
JS PY JS PY JS PY JS PY JS PY
GPT-4o Direct $55.9$ $64.5$ $31.3\%$ $28.4\%$ $11.6\%$ $29.1\%$ $12.7\%$ $24.3\%$ $14.2\%$ $26.6\%$
CoT $94.8$ $94.4$ $12.3\%$ $8.4\%$ $7.7\%$ $7.3\%$ $3.5\%$ $7.1\%$ $1.1\%$ $0.4\%$
Qwen2.5-72B-Instruct Direct $54.1$ $54.9$ $32.8\%$ $25.2\%$ $17.0\%$ $12.7\%$ $19.3\%$ $26.2\%$ $45.9\%$ $40.6\%$
CoT $86.0$ $90.4$ $19.7\%$ $18.9\%$ $3.2\%$ $5.5\%$ $7.8\%$ $15.7\%$ $10.4\%$ $8.5\%$
Cross-language consistency of perturbation effects.
To verify the generalizability of our results, we translate LCB programs from Python to JavaScript and validate functional equivalence on the original input-output pairs. As shown in Table 7, both GPT-4o and Qwen2.5-72B exhibit consistent degradation across languages. Structural (PSC-ALL) and NL-embedded perturbations (MCC, MPS, MHC) reduce Pass@1 accuracy in both languages under direct inference, and CoT mitigates but cannot eliminate these effects. Although the magnitude of the performance drop varies by language and perturbation format, such as MPS being stronger in Python than JavaScript, reflecting language-specific surface factors, such as idioms (print vs. console.log), that modulate how the model perceives cues. Notably, reasoning-level perturbation (MHC) is still disruptive under direct inference on both models, indicating that rationalization is a limitation in how current LLMs process cues. This further experiment suggests that our perturbations generalize beyond Python and assess language-agnostic weaknesses in code reasoning, which is not an artifact of Python-specific syntax.
4.3 Defenses and Limitations